OpenStudy (anonymous):

Write an equation of the line that passes through the point (1,2) and has the slope, m = 2.

OpenStudy (anonymous):

eq of the line through (x1,y1) and having slope m is y-y1=m(x-x1) hope that helps

OpenStudy (anonymous):

is it y=2x+3

OpenStudy (anonymous):

recheck your multiplication

OpenStudy (anonymous):

@razor99

OpenStudy (anonymous):

y-2 = 2(x-1) y-2 =2x -2 or y=2x

OpenStudy (anonymous):

ok thks a lot @matricked 4 everything :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sahrya ur very hot

OpenStudy (anonymous):

ummm thks

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lot of folks tell me that don know y lol

OpenStudy (anonymous):

no